OpenStudy (anonymous):

If f(x) = 5x – 3 and g(x) = 4x + 7, what is f(g(x)) ?

OpenStudy (cwrw238):

replace the x in f(x) with g(x) ( i e 4x + 7)

OpenStudy (anonymous):

5(4x+7)-3 u just put the g into the f

OpenStudy (anonymous):

so just distribute and solve for x

OpenStudy (cwrw238):

solve for x!!??

OpenStudy (anonymous):

i have no idea this shyt is so flutterin' confusing lol

OpenStudy (anonymous):

i didn't mean to say flutterin

OpenStudy (jiteshmeghwal9):

First put the value of g(x) in f(g(x)) as\[f(g(x))=f(4x+7)\]now put this value of f(x) in the given equation of f(x) as\[f(4x+7)=5(4x+7)-3\]now solve this expression & gt ur answer:)

OpenStudy (jiteshmeghwal9):

mean solve 5(4x+7)-3

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

i didn't mean to say flutterin

OpenStudy (anonymous):

i have no idea this shyt is so flutterin' confusing lol

OpenStudy (jiteshmeghwal9):

\[\LARGE{lol:)}\]